OpenStudy (anonymous):

If 2 point charges are moved from 1mm apart to 5 mm apart, the force between them will be reduced by A. 1/5 B. 1/25 C. 1/10 D. 1/2.5

OpenStudy (irishboy123):

\(F_A = k \dfrac{q_1q_1}{r^2}\) \(F_B = k \dfrac{q_1q_1}{(5r)^2}\) **reduced by** is also plain bad termonilogy :-)) \(\Delta F \%= \dfrac{F_B - F_A}{F_B} \) etc etc etc

OpenStudy (shamim):

U know F is inversely proportional to r^2

OpenStudy (shamim):

First time r=1mm

OpenStudy (shamim):

Second time r=5mm

OpenStudy (shamim):

So second time F will b reduced

OpenStudy (shamim):

Fb/Fa=?